# Evaluation and Facilitation of Online Discussions in the LLM Era: A Survey

### v1

@@ 5.4 Engagement and Impact
Engagement refers to the level of interest and participation in a discussion. It can be assessed by measuring response relevance, balance between questions and statements, interaction flow, and sometimes discussion length Adomavicius (2021). Zhang et al. (2024) document that engagement assignment is a challenge task for LLMs.
@@
Empirical literature has primarily investigated factors influencing persuasion, including linguistic strategies (e.g., word matching) and interaction dynamics (e.g., back-and-forth engagement Tan et al. (2016)). Researchers have also examined the semantic types of argumentative components (premises and claims) , such as ethos-based appeals and interpretationsHidey et al. (2017). Additionally, studies have explored dynamic factors related to topics and discourse Zeng et al. (2020) and have developed models aimed at accurately predicting persuasion. However, Ziems et al. (2024) highlight the poor persuasiveness annotation capability of LLMs in online argumentative discussions.
@@
Diversity improves discussion quality by introducing varied perspectives and experiences Zhang et al. (2024). Informativeness refers to the relevance and value of information shared in a discussion. Zhang et al. (2024) document that LLMs struggle to assess diversity and informativeness.
@@ 6.1 When to Intervene
Picking the right moment to intervene is a crucial part of effective facilitator strategies. If a facilitator does not intervene when they should have, there is a risk of significant escalation, while intervening when unnecessary can increase toxicity Schaffner et al. (2024); Trujillo and Cresci (2022); Schluger et al. (2022); Cresci et al. (2022). It is imperative then, that a facilitator is able to recognize subtle cues hinting towards escalation, in order to defuse the situation (something that even experienced humans are not confident to reliably do Schluger et al. (2022)). The NLP task of ‘Conversational Forecasting’ may contribute towards this direction. Given a conversation up to a point, a model attempts to predict if an event will occur in the future in that conversation. In our case, this event would be a facilitation intervention Schluger et al. (2022). One way to solve this problem models it as a binary classification task (see ‘Conversation Derailment’ datasets, §8). Traditional ML models can perform better than baselines on this task, although their performance varies Falk et al. (2021); Park et al. (2012); Falk et al. (2024); Schluger et al. (2022).
@@ 6.2 How to Intervene
There is currently no standard, agreed-upon taxonomy for facilitator interventions. Lim et al. (2011) propose a taxonomy that focuses on discussion facilitation, excluding, however, disciplinary or administrative actions, which are common in online discussions. Park et al. (2012) propose another taxonomy consisting of seven moderator functions, ranging from policing the discussion to solving technical issues, each of which corresponds to several possible intervention types. These functions roughly correlate with the volunteer moderator roles found in Seering (2020). More practical approaches can be found in facilitator manuals eRulemaking Initiative (2017); MIT Center for Constructive Communication (2024) and books White et al. (2024).
Facilitators often have to decide what form of coercive measure they will take to make sure the conversation remains healthy, without having to intervene repeatedly. Human interventions typically use an unofficial ‘escalation ladder’, where the facilitator will progressively move from standard facilitation tactics to threatening, and finally disciplinary action Seering (2020). Disciplinary action should be used as a last resort, since ‘conversational moderation’ Cho et al. (2024) (where a facilitator first converses with the offender) has proven effective, and is actively encouraged in some facilitator guidelines The Commons (2025). Indeed, it is typically not the first choice of a facilitator Schluger et al. (2022)
That said, there are also ‘softer’ kinds of interventions that facilitators frequently use, such as: setting and informing users about rules Schluger et al. (2022); Seering (2020), welcoming new users Schluger et al. (2022), summarizing key points Small et al. (2023); Falk et al. (2024), balancing participation Kim et al. (2021); Fishkin et al. (2018) and aiding users in structuring their speech Tsai et al. (2024); Falk et al. (2024).
@@ 6.3 Personalized Interventions
Finally, it is worth stressing that intervention strategies should not be applied en masse, without considering the characteristics of each individual. Traditionally, massive application (or threatening) of disciplinary action has led to adverse effects community- and platform-wide Trujillo and Cresci (2022); Falk et al. (2021) and the creation of echo-chambers Cho et al. (2024). There are also calls for research to move away from one-size-fits-all approaches and instead move towards personalized interventions Cresci et al. (2022). Human facilitators are often able to personalize interventions per individual Schluger et al. (2022), and we hypothesize that LLMs can also do so to some extent.
@@ 7 Towards LLM-based facilitation
Until recently, ML models used as facilitation agents were confined to either performing menial tasks, such as pasting automated messages Seering (2020); Schluger et al. (2022), suggesting facilitation actions (e.g., rejecting posts), possibly via human-in-the-loop frameworks Fishkin et al. (2018); Gelauff et al. (2023), identifying possibly escalatory comments Schluger et al. (2022), or employing pre-programmed facilitative tactics (such as Kim et al. (2021)), where the model produces automated messages encouraging participation). However, ML-based and rule-based facilitation are not effective enough to meet the high demands of most platforms Seering (2020); Schaffner et al. (2024).
Advances in LLMs enable the development of facilitation agents that more actively engage in discussions. These agents can warn users about policy violations Kumar et al. (2024), suggest rephrasings to improve tone or persuasiveness Bose et al. (2023), monitor turn-taking Schroeder et al. (2024), and summarize or visualize key discussion points Small et al. (2023). They can also assist in drafting group statements that reflect diverse viewpoints Tessler et al. (2024). A brief, non-exhaustive summary of the capabilities of ML models, LLMs, and humans can be found in Figure 2.
Figure 2: Current capabilities of ML, LLM, and human facilitation. From left to right: new tasks are possible, but the cost of solving them rises proportionally. In-between actors, are tasks that are solved suboptimally by the previous actor; e.g., ML systems can be used for discussion evaluation, but are not particularly effective.
@@ 7.1 Administrating the discussion
LLMs are well positioned to tackle a wealth of ‘administrative’ facilitation tasks, which aim to organize a discussion. For instance, facilitators typically periodically summarize the viewpoints of the participants and seek their confirmation in order to both make them feel understood, and present their views to other participants. This iterative summarization is a task which LLMs may be well-equipped to handle Small et al. (2023); Burton et al. (2024). According to Jin et al. (2024), LLMs bring significant advantages over traditional methods, “notably in the quality and flexibility of the generated texts and the prompting paradigm to alleviate the cost of training deep models”. Specifically on discussion summarization, however, Feng and Qin (2022) suggest that the challenges are profound, as discussions contain multiple participants, topic drifts, multiple co-references, diverse interactive signals, and diverse domain terminologies.
In some deliberative contexts, facilitators are also encouraged to begin a discussion with their own opinion Small et al. (2023), although others disagree MIT Center for Constructive Communication (2024). This is a task LLMs can also handle, albeit less convincingly than current Information Retrieval (IR) approaches Karadzhov et al. (2021).
Finally, LLMs can help marginalized groups in discussions by offering translations of the discussion in their native languages, and by helping them phrase their opinions with proper grammar and syntax Tsai et al. (2024); Burton et al. (2024).
@@ 7.2 Evolving traditional automation models
LLMs have been proven to be adept at NLP tasks such as the detection of hate speech Shi et al. (2024), toxicity Kang and Qian (2024); Wang and Chang (2022), and misinformation (fake news) Kang and Qian (2024); Wang and Chang (2022). These abilities make LLMs usable as drop-in replacements for traditional ML models for these tasks. They also suggest that conversational LLM facilitation agents may be able to identify, and dynamically adapt to such phenomena properly.
7.3 Fully Automatic LLM-based facilitation
There are indications that LLM chatbots can be used as facilitators in the fullest capacity of the role. LLMs are able to predict optimal facilitation tactics Schroeder et al. (2024), like traditional ML models Al-Khatib et al. (2018). Furthermore, they have proven capable of developing and executing social strategies in other tasks, e.g., negotation games, LLM interactions Abdelnabi et al. (2024); Cheng et al. (2024); Martinenghi et al. (2024). Given that relatively simple ML chatbots Kim et al. (2021), which do not leverage generative text capabilities, succeed at improving discussions, many expect LLM-based facilitation to be a promising solution to the well-known bottleneck of human facilitation Small et al. (2023); Seering (2020); Burton et al. (2024); Schroeder et al. (2024), with some applications already showing promise Cho et al. (2024).
@@ 8 Facilitation Datasets
In this section, we provide an overview of the most prominent datasets for online moderation, considering their sizes and their relevance to core facilitation tasks. We propose the following new taxonomy of facilitation datasets: Conversation Derailment datasets, where the task is to predict when a conversation escalates and requires facilitator intervention; User Tactics datasets, which concern how users position themselves during the discussion; and Facilitator Interventions datasets. Some datasets contain information that can be used in multiple tasks. An overview of the surveyed datasets and their categories in our taxonomy can be found in Table 1.
Name Task Size Content Wikipedia Disputes De Kock and Vlachos (2021) Conversation Derailment User Tactics 7,42574257,4257 , 425 D Includes annotations for several ‘dispute tactics.’ WikiConv Hua et al. (2018) User Tactics 91,000,0009100000091,000,00091 , 000 , 000 D Includes moderation meta-data such as comment edits and deletions. Webis-WikiDebate-18 Al-Khatib et al. (2018) User Tactics 6,000,00060000006,000,0006 , 000 , 000 D Graph-based, includes data annotated for argumentation strategies. Conversations Gone Awry Zhang et al. (2018b) Conversation Derailment User Tactics 4,18841884,1884 , 188 D Predicts derailment by analyzing rhetorical tactics, human-annotated. Chang and Danescu-Niculescu-Mizil (2019) (1) Conversation Derailment 4,18841884,1884 , 188 D Extends the ‘Conversations Gone Awry’ dataset. Chang and Danescu-Niculescu-Mizil (2019) (2) Conversation Derailment 6,84268426,8426 , 842 D Based on the r/ChangeMyView subreddit. Park et al. (2012) Conversation Derailment Facilitator Interventions 1,67816781,6781 , 678 C Comprised of 4 datasets. Includes 19 intervention types belonging to 7 moderator roles. RegulationRoom Falk et al. (2021) Conversation Derailment 3,00030003,0003 , 000 C Extends the dataset of Park et al. (2012). DeliData Karadzhov et al. (2021) User Tactics Facilitator Interventions $500500$ D Group discussions, includes task-oriented quality measure which may be used to approximate discussion quality. Wiki-Tactics De Kock et al. (2022) Facilitator Interventions $213213$ D Based on Wikipedia Disputes, includes moderation action metadata such as comment edits and deletions. UMOD Falk et al. (2024) Facilitator Interventions 2,00020002,0002 , 000 C Based on the r/ChangeMyView subreddit, annotated for facilitation tactics and AQ. Fora Schroeder et al. (2024) Facilitator Interventions $262262$ D Original dataset revolving around experience-sharing, annotated for facilitation tactics.
Table 1: Overview of reviewed datasets. Unnamed datasets are referred to by the names of the authors. The size reflects the number of annotated conversations, disregarding unlabeled data. In this table, D refers to the number of discussions, while C, to the number of individual comments.
@@ 9 LLM Discussion Facilitation Roadmap
LLMs can serve as automated discussion quality annotators. Are these annotators infallible? At the moment, the answer is no. There are discussion quality aspects, with specific characteristics, that LLMs cannot annotate in a highly accurate way. Particularly, aspects characterized by highly subjective and complicated expert taxonomies (e.g., empathy), as well as specific AQ dimensions requiring the integration of social norms (e.g., diversity; see §5). However, even human annotations tend to be polarized for such subjective quality dimensions (e.g., toxicity) due to sociodemographic background effects and personal biases Beck et al. (2024); Sap et al. (2020).
Unlike traditional automated (self-)supervised techniques that require costly hardware and extensive training data, prompted LLMs offer a more affordable alternative for discussion quality annotation without the need for expensive training examples. Yet, as any tool, LLMs should be used with caution. For each quality aspect of the discussion, researchers must choose the proper models (e.g., open/close source, the size, trained or not with reinforcement learning, fine-tuned or not on specialized data). Moreover, prompts specifically engineered for each quality aspect may also be adopted.
Future research should consider practical evaluation frameworks for both the meta-evaluation of discussion quality metrics—assessing their effectiveness, consistency, bias, robustness, and real-world applicability—and for moderation and facilitation interventions. Evaluating such frameworks and facilitation strategies requires rigorous testing in controlled environments. However, real-world evaluations can be costly, time-consuming, and difficult to replicate, making alternative approaches necessary. To this end, synthetic experiments made exclusively with LLM actors may be of great help, although future work will have to prove that synthetic experiments are a reliable proxy for human behavior to a certain degree Hewitt et al. (2024); Park et al. (2023, 2022, 2024).
@@
Intervention types should be adapted to the different legal frameworks and rules of each community/platform. While there are exhaustive surveys on the topic, such as that of Schaffner et al. (2024), there is yet no methodology to train human or synthetic facilitators according to these factors. We posit that experiments using exclusively LLM user/facilitator-agents are necessary to sustainably test new facilitation strategies and interventions, as applied in other NLP tasks Ulmer et al. (2024); P.Cheng et al. (2024); Park et al. (2022, 2023). Finally, the datasets presented in Table 1 can be used to train and assess LLM facilitators in the future, as well as to generate additional data—similar to the existing ones but with controlled modifications—to stress-test various facilitation training settings.

### v2

@@ 4.4 Engagement and Impact
Engagement
Engagement is desirable in online discussion platforms as it combines interest and participation Lambert et al. (2024); Park et al. (2012). It is proxied by measures like reciprocity Graham and Witschge (2003); Stromer-Galley (2007); Zhang et al. (2018), number of comments posted by each user Avalle et al. (2024), discussion length Adomavicius (2021); Avalle et al. (2024), while Ferron et al. (2023) define subdimensions such as response diversity, interestingness, and specificity.
@@
Empirical literature has primarily examined factors influencing persuasion that align with other categories in our taxonomy, such as linguistic markers (§4.1) and turn-taking (§4.2)Tan et al. (2016). Considering this connection, persuasion is not only an indicator of argument quality, but may also serve as a proxy for identifying additional markers signaling whether facilitator intervention is needed.
@@
Diversity in online discussions refers to the presence of varied perspectives, backgrounds, and experiences, which can enrich conversations by fostering constructive exchanges Irani et al. (2024); Zhang et al. (2024). To prevent echo chambers and promote inclusivity, facilitators can use diversity measures to encourage opinion diversity Anastasiou et al. (2023), encouraging users to explore a broad range of perspectives on a given issue Kim et al. (2021). Informativeness refers to the relevance and value of information shared in a discussion and is considered a building stock of prosociality, an attribute that facilitation trys to foster in online discussion platforms Lambert et al. (2024).
4.5 LLM Approaches to Discussion Quality
LLMs can significantly aid in evaluating discussion quality, performing on par with humans in annotating argument structure, coherence, and flow across tasks like argument mining and synthesis Mirzakhmedova et al. (2024); Rescala et al. (2024); Wang et al. (2023); Chen et al. (2024a); Irani et al. (2024); Anastasiou and De Liddo (2024); Zhang et al. (2024); Mendonca et al. (2024); Zhang et al. (2023). LLMs can also reliably label dialogue acts Ziems et al. (2024); Cimino et al. (2024); Martinenghi et al. (2024); Schroeder et al. (2024), as well politeness, power, disagreement, and toxicity Zhou et al. (2024); Ziems et al. (2024). However, they struggle with tasks involving social norms (such as irony and humor) and show limited accuracy in sentiment and engagement detection Hu et al. (2023); Sravanthi et al. (2024); Furniturewala and Jaidka (2024); Xu and Jiang (2024). While LLMs show promise in measuring controversy and persuasion, performance drops at the discussion level, especially when aiming to measure diversity, informativeness, and generally when social and pragmatic understanding is necessary Ziems et al. (2024); Avalle et al. (2024); Lawrence and Reed (2020).
@@ 5.1 When to Intervene
Picking the right moment to intervene is a crucial part of effective facilitation strategies. If a facilitator does not intervene when they should have, there is a risk of significant escalation, while intervening when unnecessary can increase toxicity Schaffner et al. (2024); Trujillo and Cresci (2022); Schluger et al. (2022); Cresci et al. (2022). It is imperative then (also considering the evaluation dimensions discussed in §4), for a facilitator to be able to recognize subtle cues hinting towards escalation, in order to defuse the situation, something that even experienced human facilitators are not confident to reliably do Schluger et al. (2022). The NLP task of ‘Conversational Forecasting’ may contribute towards this direction. Given a conversation up to a point, a model attempts to predict if an event will occur in the future in that conversation. In our case, this is where a facilitator would intervene Schluger et al. (2022). Traditional ML models can perform well on this task, although their performance varies Falk et al. (2021); Park et al. (2012); Falk et al. (2024); Schluger et al. (2022).
@@ 5.2 How to Intervene
There is currently no agreed-upon taxonomy for facilitator interventions. Lim et al. (2011) propose a taxonomy that focuses on discussion facilitation, excluding, however, disciplinary or administrative actions, which are common in online discussions. Park et al. (2012) propose another taxonomy consisting of seven moderator functions, ranging from policing the discussion to solving technical issues. These functions roughly correlate with the volunteer moderator roles, as described by Seering (2020). More practical approaches can be found in facilitator manuals eRulemaking Initiative (2017); MIT Center for Constructive Communication (2024) and books White et al. (2024).
Facilitators often have to decide what form of coercive measure to take to make sure the conversation remains healthy, without having to intervene repeatedly. Human interventions typically use an unofficial ‘escalation ladder’ (Figure 1), where the facilitator will progressively move from milder facilitation tactics to threatening, and finally disciplinary action Seering (2020). ‘Conversational moderation’ Cho et al. (2024), where a facilitator first converses with the offender, has proven effective and is actively encouraged in some facilitator guidelines The Commons (2025). This is probably why disciplinary action is typically not the first choice of a facilitator Schluger et al. (2022) and why it should reasonably be used as a last resort. Softer kinds of interventions that facilitators frequently use first include: setting and informing users about rules Schluger et al. (2022); Seering (2020), welcoming new users Schluger et al. (2022), summarizing key points Small et al. (2023); Falk et al. (2024), balancing participation Kim et al. (2021); Fishkin et al. (2018), and aiding users improve their points Tsai et al. (2024); Falk et al. (2024).
@@ 5.3 Personalized Interventions
It is worth stressing that intervention strategies should not be applied en masse, without considering the characteristics of each individual. Traditionally, massive application (or threatening) of disciplinary action has led to adverse effects community- and platform-wide Trujillo and Cresci (2022); Falk et al. (2021) and the creation of echo-chambers Cho et al. (2024). There are also calls for research to move away from one-size-fits-all approaches and instead move towards personalized interventions Cresci et al. (2022). Human facilitators are often able to personalize interventions per individual Schluger et al. (2022), and we hypothesize that LLMs can also do so to some extent.
@@ 6 Towards LLM-based facilitation
Until recently, ML models used as facilitation agents were confined to either performing menial tasks, such as pasting automated messages Seering (2020); Schluger et al. (2022), suggesting facilitation actions (e.g., rejecting posts), possibly via human-in-the-loop frameworks Fishkin et al. (2018); Gelauff et al. (2023), identifying possibly escalatory comments Schluger et al. (2022), or employing pre-programmed facilitative tactics, as in the work of Kim et al. (2021), where the model produces automated messages encouraging participation. However, older ML-based and rule-based facilitation are not effective enough to meet the high demands of most platforms Seering (2020); Schaffner et al. (2024).
Advances in LLMs enable the development of facilitation agents that more actively engage in discussions. These agents can warn users about policy violations Kumar et al. (2024), suggest rephrasings to improve tone or persuasiveness Bose et al. (2023), monitor turn-taking Schroeder et al. (2024), and summarize or visualize key discussion points Small et al. (2023). They can also assist in drafting group statements that reflect diverse viewpoints Tessler et al. (2024). A brief, non-exhaustive summary of the capabilities of simpler ML models, LLMs, and humans can be found in Figure 2.
Figure 2: Capabilities of simpler ML, LLM, and human facilitation. Task complexity and cost increase from left to right. Intermediate tasks are handled suboptimally by the preceding method.
@@ 6.1 Administrating the Discussion
LLMs are able to tackle a variety of ‘administrative’ facilitation tasks that help structure discussions. For example, facilitators often summarize the views of the participants, seek confirmation of understanding, and share perspectives. This iterative summarization is a task LLMs may handle effectively Small et al. (2023); Burton et al. (2024). However, Feng and Qin (2022) point out some challenges such as discussions with multiple participants, topic drifts, multiple co-references, diverse interactive signals, and diverse domain terminologies. Still, according to Jin et al. (2024), LLMs bring significant advantages over conventional ML methods, “notably in the quality and flexibility of the generated texts and the prompting paradigm to alleviate the cost of training deep models”.
In some deliberative contexts, facilitators are also encouraged to begin a discussion with their own opinion Small et al. (2023), although others disagree MIT Center for Constructive Communication (2024). This is a task LLMs can also handle, albeit less convincingly than Information Retrieval (IR) approaches Karadzhov et al. (2021).
Finally, LLMs can help marginalized groups in discussions by offering translations of the discussions in their native languages, and by helping them phrase their opinions with proper grammar and syntax Tsai et al. (2024); Burton et al. (2024). This can directly improve discussions by increasing their diversity (Section 4.4).
@@ 6.2 Evolving Traditional Automation Models
LLMs have been proven to be adept at NLP tasks such as the detection of hate speech Shi et al. (2024), toxicity Kang and Qian (2024); Wang and Chang (2022), and misinformation Kang and Qian (2024); Wang and Chang (2022). These abilities make LLMs usable as drop-in replacements for traditional ML models for these tasks, suggesting that conversational LLM facilitation agents may be able to identify, and dynamically adapt to such phenomena properly. We note however that LLMs are much more expensive and less scalable than their simpler ML counterparts. Furthermore, LLM annotation has its own challenges: LLM survey responses Jansen et al. (2023); Bisbee et al. (2024); Neumann et al. (2025) and annotations Gligori’c et al. (2024) are generally unreliable and surface-level. Non-deterministic behavior is also common in LLMs Atil et al. (2025), but also particularly in closed-source models Bisbee et al. (2024) on which a lot of research on LLM annotation hinges.
6.3 Fully Automated LLM-based Facilitation
There are indications that LLMs can be used as facilitators in the fullest capacity of the role. LLMs are able to predict optimal facilitation tactics Schroeder et al. (2024), like traditional ML models Al-Khatib et al. (2018). Furthermore, they have proven capable of developing and executing social strategies in other tasks, e.g., negotation games, LLM interactions Abdelnabi et al. (2024); Cheng et al. (2024a); Martinenghi et al. (2024). Given that relatively simple ML chatbots, which do not leverage generative text capabilities, have been reported to improve discussions Kim et al. (2021), many expect LLM-based facilitation to be a promising solution to the well-known bottleneck of human facilitation Small et al. (2023); Seering (2020); Burton et al. (2024); Schroeder et al. (2024). Notably, Cho et al. (2024) successfully use LLM facilitators with prompts based on Cognitive Behavioral Therapy to moderate a live discussion with human participants. Their work shows that LLM facilitators can adapt in their instructions to users, although they cannot by themselves affect the discussion with regard to cooperation and mutual respect between the participants.
Nevertheless, LLMs have inherent limitations that make them worse than humans in most social tasks (Figure 2) Rossi et al. (2024). While human facilitators are encouraged to be neutral White et al. (2024); eRulemaking Initiative (2017), numerous studies point to biases in sociodemographic, statistical, and political terms in LLMs R.Anthis et al. (2025); Hewitt et al. (2024); Rossi et al. (2024), which can be exacerbated during the course of a discussion Taubenfeld et al. (2024).
@@ 7 Facilitation Datasets
In this section, we provide an overview of the most prominent datasets for online facilitation, considering their sizes and their relevance to core facilitation tasks. We propose the following new taxonomy of facilitation datasets: Conversation Derailment datasets, where the task is to predict when a conversation escalates, therefore requiring facilitator intervention; and Facilitator Interventions datasets, which include comments by facilitators in active discussions, sometimes annotated with the tactics employed. Some datasets contain information that can be used in multiple tasks. An overview of the surveyed datasets and their categories in our taxonomy can be found in Table 1.
Name Task Size Content Wikipedia Disputes De Kock and Vlachos (2021) Conversation Derailment 7,42574257,4257 , 425 D Includes annotations for several ‘dispute tactics.’ WikiConv Hua et al. (2018) Facilitator Interventions 91,000,0009100000091,000,00091 , 000 , 000 D Includes moderation meta-data such as comment edits and deletions. Conversations Gone Awry Zhang et al. (2018) Conversation Derailment 4,18841884,1884 , 188 D Predicts derailment by analyzing rhetorical tactics, human-annotated. Chang and Danescu-Niculescu-Mizil (2019) (1) Conversation Derailment 4,18841884,1884 , 188 D Extends the ‘Conversations Gone Awry’ dataset. Chang and Danescu-Niculescu-Mizil (2019) (2) Conversation Derailment 6,84268426,8426 , 842 D Based on the r/ChangeMyView subreddit. Park et al. (2012) Conversation Derailment Facilitator Interventions 1,67816781,6781 , 678 C Comprised of 4 datasets. Includes 19 intervention types belonging to 7 moderator roles. RegulationRoom Falk et al. (2021) Conversation Derailment 3,00030003,0003 , 000 C Extends the dataset of Park et al. (2012). DeliData Karadzhov et al. (2021) Facilitator Interventions $500500$ D Group discussions, includes task-oriented quality measure which may be used to approximate discussion quality. Wiki-Tactics De Kock et al. (2022) Facilitator Interventions $213213$ D Based on Wikipedia Disputes, includes moderation action metadata such as comment edits and deletions. UMOD Falk et al. (2024) Facilitator Interventions 2,00020002,0002 , 000 C Based on the r/ChangeMyView subreddit, annotated for facilitation tactics and AQ. Fora Schroeder et al. (2024) Facilitator Interventions $262262$ D Original dataset revolving around experience-sharing, annotated for facilitation tactics.
Table 1: Overview of reviewed datasets. Unnamed datasets are referred to by the names of the authors only. The size reflects the number of annotated conversations, disregarding unlabeled data. D indicates the number of discussions. C indicates the number of individual comments or dialogue turns.
@@ 8 LLM Discussion Facilitation Roadmap
LLMs can serve as automated discussion quality annotators (§4). Are these annotators infallible? Not yet. Certain dimensions, especially those that are highly subjective (e.g., pragmatic understanding), remain challenging for LLMs to annotate accurately. But we must take into account that even human annotations tend to be polarized for such subjective quality dimensions Argyle et al. (2023), largely due to sociodemographic background effects and personal biases Beck et al. (2024); Sap et al. (2020).
On the other hand, prompted LLMs offer a more scalable and cost-effective alternative for annotating discussion quality compared to human annotation and traditional (or self-) supervised training on large annotated datasets. Using LLMs for annotation, however, requires careful model selection considering whether models are open or closed source, model size, model alignment, as well as prompt selection, and (if applicable) fine-tuning requirements. These choices should be tailored to the specific quality dimension being evaluated.
@@
Intervention types should be adapted to the different legal frameworks, rules, and social norms of each community/platform. While there are exhaustive surveys on intervention types and policies, such as that of Schaffner et al. (2024), there is yet no methodology to train human or artificial facilitators according to these factors. We posit that experiments using exclusively LLM user/facilitator-agents are necessary to sustainably test new facilitation strategies and interventions per community and platform, as in other NLP tasks that involve LLM-generated conversation Ulmer et al. (2024); Cheng et al. (2024b); Park et al. (2022, 2023), before testing the resulting facilitators in costly experiments with human participants. Finally, the datasets presented in Table 1 can be used to train and assess LLM facilitators in the future, as well as to generate additional data—similar to the existing ones, but with controlled modifications—to stress-test various facilitators in particular settings (e.g., predicting or recovering from a conversation derailment).